OpenStudy (cherrie):

An artifact was found and tested for its carbon-14 content. If 82% of the original carbon-14 was still present, what is its probable age (to the nearest 100 years)? Use that carbon-14 has a half-life of 5,730 years.

OpenStudy (anonymous):

The basic formula for this kind of problem is \[\large N = No\ e^{-kt}\]

OpenStudy (anonymous):

\[\large .82=\left(\frac{1}{2}\right)^{\frac{t}{5730}}\] solve for \(t\)

OpenStudy (anonymous):

Another useful thing is that if the half-life is T T k = ln(2)

OpenStudy (anonymous):

You want to work through that one, I think?

OpenStudy (anonymous):

you have a choice, you can use @telliott99 method, find \(k\) and then solve, or you can solve the equation i wrote using the change of base formula that method gives \[t=5730\times \frac{\ln(.83)}{\ln(.5)}\]

OpenStudy (anonymous):

@satellite73 sure you want that 1/2 in the base

OpenStudy (anonymous):

Is that part of your change of base, I have to work through that ..